OpenStudy (4meisu):

Find the equation of the line containing (4.-3) that is parallel to the line 6x-4y=1

OpenStudy (anonymous):

parallel line means same slope in order to find the slope of : 6x-4y=1 we will change it to slope,y intercept form : 4y = 6x - 1 y = 1.5x -0.25 so the slope is 1.5 now we will use line equation : y= y1 + m(x-x1) and plug the slope and the point : y = -3 +1.5(x-4) y = -3 + 1.5x - 6 y = 1.5x - 9